OpenStudy (anonymous):

Assume a jar has 7 red marbles and 3 black marbles. Draw out 3 marbles, with and without replacement. Find the requested probabilities. a) P( 3 black marbles) with replacement, without replacement b) P( 1 red marble and 2 black marbles) with replacement, without replacement. c) P( red on the first draw and black on the second draw, and black on the third draw) with replacement, without replacement.

OpenStudy (anonymous):

3 black marbles with replacement \[\frac{3}{10}\times \frac{3}{10}\times \frac{3}{10}=\left(\frac{3}{10}\right)^3\] without replacement \[\frac{3}{10}\times \frac{2}{9}\times \frac{1}{8}\] others are similar

OpenStudy (anonymous):

That same technique doesn't work with the others :(

OpenStudy (anonymous):

I figured out "c"! But everything I try for "b" doesn't work...